# Time Series Forecasting with LLMs: Understanding and Enhancing Model Capabilities

### v1

@@ Abstract
Large language models (LLMs) have been applied in many fields with rapid development in recent years. As a classic machine learning task, time series forecasting has recently received a boost from LLMs. However, there is a research gap in the LLMs’ preferences in this field. In this paper, by comparing LLMs with traditional models, many properties of LLMs in time series prediction are found. For example, our study shows that LLMs excel in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity. We explain our findings through designing prompts to require LLMs to tell the period of the datasets. In addition, the input strategy is investigated, and it is found that incorporating external knowledge and adopting natural language paraphrases positively affects the predictive performance of LLMs for time series. Overall, this study contributes to insight into the advantages and limitations of LLMs in time series forecasting under different conditions.
@@ 1 Introduction
Recently, large language models (LLMs) have been widely used, achieving promising performance across various domains, such as health management, customer analysis, and text feature mining Peng et al. (2023); Ledro et al. (2022); Huang et al. (2023). Time series forecasting requires extrapolation from sequential observations. Language models, designed to discern intricate concepts within temporally correlated sequences, intuitively appear well-suited for this task. Hence, LLMs demonstrate proficiency in the domain of time series forecasting Gruver et al. (2023); Rasul et al. (2023); Sun et al. (2023).
Currently, the application of LLMs for time series prediction is at a nascent stage, with the boundaries of this research area remaining ill-defined. The utilization of proprietary large language models in this domain lacks a clear, established methodology. Our objective is to shed light on this emerging field, offering valuable insights and guidance for future research endeavors. However, the LLMs’ preferences for the input time series remain unexplored, such as their proficiency in forecasting both seasonal and trend data. We believe that studying this area clearly will greatly improve the performance of LLM in time series forecasting.
@@
To fill this research gap, in this paper, we focus on the question of what are LLMs’ preferences for the input time series in time series forecasting. To answer this, we conduct experiments on both real and synthesized datasets. Our observations reveal that LLMs perform better on time series characterized by higher trend or seasonal strengths. To further discern the LLMs’ preferences for the specific segments of the input data, we design counterfactual experiments involving systematic permutations of input sequences. We find that LLMs are sensitive to the segment of input sequences proximate to the output.
Naturally, we are interested in the question: Why do LLMs forecast well on datasets with higher trend or seasonal strengths? To solve this, we design prompts that require LLMs to tell the period of the datasets. Through experiments, we let the large language model tell the period of the dataset several times and take the median. We found that the large language model can accurately pinpoint the periodicity of a dataset. This can explain why large language models can predict data sets with high trends or seasonal intensities well since they already learned this kind of knowledge.
@@
In light of these findings, our focus lies on how to leverage these insights to further improve model performance. To address this, we propose two simple techniques to enhance model performance: incorporating external human knowledge and converting numerical sequences into natural language counterparts. Incorporating supplementary information enables large language models to more effectively grasp the periodic nature of time series data, moving beyond a mere emphasis on the tail of the time series. Transforming numerical data into a natural language format enhances the model’s ability to comprehend and reason, also serving as a beneficial approach. Both approaches improve model performance and contribute to our understanding of LLMs in time series forecasting. The workflow is illustrated in Figure 1.
@@
We investigate the LLMs’ preferences for the input sequences in time series forecasting. Based on our analysis, we have observed that LLMs are more effective in datasets that have clear periods and trends. A surprising finding is that LLMs may generate output sequences that exhibit higher trend and seasonal strengths. Besides, our results reveal a limitation in LLMs’ performance when confronted with multi-period time series datasets. It indicates that LLMs face challenges in recognizing the distinct periods inherent in such datasets. To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian to the original time series to create counterfactual instances. Our results indicate that large language models are sensitive to the segment of input sequences proximate to the output. All these findings suggest a potential area for improvement in LLMs’ adaptability to complex temporal structures.
Through prompt design and the use of in-context Learning, we have verified that LLMs make better predictions on datasets that exhibit strong trends or seasonality, showing that these models can precisely identify dataset periodicity.
We propose two simple techniques to improve model performance and find that incorporating external knowledge in prompts and paraphrasing natural language positively affects the performance of LLMs in time series forecasting.
@@ 2.1 Large Language Model
We use LLMs as a zero-shot learner for time series forecasting by treating numerical values as text sequences. The success of LLMs in time series forecasting can significantly depend on correct pre-processing and handling of the data Gruver et al. (2023). We followed their approach and this process involves a few crucial steps.
In the pre-processing phase for time series forecasting with LLMs, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing. For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to "1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0", introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space. Tokenization is equally pivotal, shaping the model’s pattern recognition capabilities. Unlike traditional methods like byte-pair encoding (BPE) Hugging Face (2023), which may disrupt numerical coherence, spacing digits ensures individual tokenization, enhancing pattern discernment.
Moreover, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so a specific percentile aligns to 1, facilitating the model’s exposure to varying digit counts and supporting the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
@@ 2.2 Time Series Forecasting
Here, represents the -step estimation given the previous -step values . denotes the trained parameters from the model , and denotes the prompt or any other information used for the inference. This paper predominantly focuses on single-variate time series forecasting.
Motivated by the interpretability requirements in real-world scenarios, time series can often be decomposed into the trend component, the seasonal component, and the residual component through the addictive model Cleveland et al. (1990). The trend component captures the hidden long-term changes in the data, such as the linear or exponential pattern. The seasonal component captures the repeating variation in the data, and the residual component captures the remaining variation in the data after the removal of trend and seasonal components. This decomposition offers a method to quantify the properties of the time series, which is detailed in section 3.2.
@@ 3 What are LLMs’ preferences in time series forecasting?
To explore the preference in the large language model, we first quantify the properties of the input time series to investigate the LLMs’ preferences for time series. Then, to further emphasize our findings, we evaluate the importance of different segments of the input sequence by adding Gaussian noise to the original time series.
@@ 3.1 Analyzing Method
To understand the preferences of the LLMs in time series forecasting, we compare LLMTime using various foundational models like GPT-4 and GPT-3.5-turbo traditional methods on various datasets. We also design experiments on synthesized datasets to validate our findings and analyze the impact of the multiple periods. To quantify the LLMs’ preferences towards time series, following Wang et al. (2006), we define the strength of the trend and the seasonality as follows:
where , and denote the trend component, the seasonal component and the residual component respectively. The presented indices serve as indicators of the strength of the trend and seasonality, providing a measure ranging up to 1. It is easy to find that a higher value indicates a stronger trend or seasonality within the time series.
To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian to the original time series to create counterfactual examples. We initiate by defining a sliding window that constitutes 10% of the total length of the time series. Given the variability in the length of time sequences, we first scale the sequence before incorporating the noise into the original sequence. This method allows us to assess the impact of individual segments and thereby infer the interpretability of the time series segments that LLMs predominantly focus on.
3.2 Preferences for the time series Experiment
In this subsection, we delve into the input sequence preferences for time series forecasting with LLMs. Our experiments are conducted on both real datasets and synthetic datasets. Throughout this subsection, we use GPT-3.5-turbo-instruct and GPT-4 for time series forecasting and measure model performance through and Mean Absolute Percentage Error (MAPE).
@@ 3.2.1 Implementation Details
Real Datasets: We conduct experiments on ten real datasets, as enumerated in Appendix A.2. we apply the Seasonal-Trend decomposition using the LOESS (STL) techniqueCleveland et al. (1990), to decompose the original time series into trend, seasonal, and residual components. Subsequently, we computed the strengths of the trend strength and seasonal strength . To further discern the LLMs’ preferences for the specific segments of the input data, we conduct the counterfactual analysis with a systematic permutation to the input time series. We initiate by defining a sliding window that constitutes 10% of the total length of the time series and add Gaussian noise into the data within this window data. Given the variability in the length of time sequences, we first scale the sequence through max-min normalization before incorporating it into the original sequence. Refer to A.2.1 for detailed information.
Synthesized Datasets: To further validate our findings, we conduct experiments on synthesized datasets. We formulate the dataset through the expression . To investigate the influence of the number of periods on model performance, we generated a dataset using the function . ranges from 0 to 20 and follows the normal distribution . Refer to A.2.1 for detailed information.
@@ 3.2.2 Key Findings
After computing the Pearson correlation coefficients (PCC), we observed a nearly strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series has a higher trend and seasonal strength. Notably, GPT-4 achieved a higher PCC compared to GPT-3.5-turbo-instruct. It may be attributed to human feedback during GPT-4 training, as individuals may be more aware of seasonal and trend data. Interestingly, there is an increase in the of the output generated by GPT-4, compared to the original test sequences. It indicates that GPT-4 tends to forecast time series with high seasonal strength, which may provide insights for further research. In the context of multi-period time series, the model performance decreases as the number of periods increases. It indicates that LLMs may have trouble recognizing the multiple periods inherent in such datasets, which is common in reality. For the counterfactual analysis, as shown in Figure 6, there is a noticeable decrease in values when Gaussian noise is added to the latter segments. Our findings reveal that LLMs are more sensitive to the end of input time series data when it acts as a time series predictor.
4 Why do LLMs forecast well on datasets with higher trend or seasonal strengths?
Our findings indicate that LLMs exhibit enhanced performance on time series data characterized by pronounced trends or significant seasonal variations. This raises the question: Why do LLMs excel in forecasting datasets with marked trend or seasonal patterns? To explore this phenomenon, we crafted prompts that necessitate LLMs to acknowledge the temporal span of the dataset.
This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct trend or seasonal attributes in-context learningvanced pattern recognition capabilities. By explicitly prompting LLMs to consider the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, thereby shedding light on the mechanisms that underpin their superior performance in such contexts.
@@ 4.1 Implementation Details
In order to explore the phenomenon that LLMs forecast well on datasets with higher trend or seasonal strengths, we design experiments to verify this phenomenon. We design prompts to let the LLMs output the predicted value after each sequence prediction. The target model of the experiment is GPT-3.5-Turbo (gpt-3.5-turbo-1106), and the main role of our prompt is to perform in-context learning and require output cycles, as detailed in the Appendix A.3. We select eight time series data sets such as AirPassengersDataset, count the period value after ten times of prediction, and compare the median of these ten results with the real period. The results are shown in Table 3.
| Dataset Name | GPT4-MAPE | GPT4-R2 | GPT3.5-MAPE | GPT3.5-R2 | Trend Strength | Seasonal Strength |
| AirPassengersDataset | ||||||
| AusBeerDataset | ||||||
| MonthlyMilkDataset | ||||||
| SunspotsDataset | ||||||
| WineDataset | ||||||
| WoolyDataset | ||||||
| IstanbulTrafficGPT | ||||||
| GasRateCO2Dataset | ||||||
| HeartRateDataset | ||||||
| TurkeyPower |
@@ 4.2 Key Findings
According to the results, we find that large language models can determine the periodicity of a dataset to some extent. Although the fluctuation of each time series prediction is large, the prediction on AirPassenger, Sunspots and Wooly datasets is relatively accurate, and the predicted values on other datasets are also close to our real cycle value. We speculate that LLMs can predict datasets with high trend or seasonal intensity well because they already possess some knowledge about the scene and content of the dataset.
5 How to leverage these insights to improve model’s performance?
Based on these findings, our focus is on how to leverage these findings to further improve model performance. Because our ultimate goal is to improve the time series prediction performance of the model. Therefore, we retain the original structure of the time series data but enhance the model’s input with prompts infused with additional knowledge. These prompts are crafted with seasonal trends and cyclical behaviours in the dataset to endow the model with a richer contextual understanding. In another way, we transform time series data into formats that resemble natural language sequences, enabling large language models to leverage their superior inference skills more effectively in the realm of time series prediction.

### v3

@@ Abstract
Large language models (LLMs) have been applied in many fields and have developed rapidly in recent years. As a classic machine learning task, time series forecasting has recently been boosted by LLMs. Recent works treat large language models as zero-shot time series reasoners without further fine-tuning, which achieves remarkable performance. However, there are some unexplored research problems when applying LLMs for time series forecasting under the zero-shot setting. For instance, the LLMs’ preferences for the input time series are less understood. In this paper, by comparing LLMs with traditional time series forecasting models, we observe many interesting properties of LLMs in the context of time series forecasting. First, our study shows that LLMs perform well in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity. This observation can be explained by the ability of LLMs to recognize the underlying period within datasets, which is supported by our experiments. In addition, the input strategy is investigated and it is found that incorporating external knowledge and adopting natural language paraphrases substantially improve the predictive performance of LLMs for time series. Overall, our study contributes insight into LLMs’ advantages and limitations in time series forecasting under different conditions.
@@ 1 Introduction
Recently, large language models (LLMs) have been widely used and have achieved promising performance across various domains, such as health management, customer analysis, and text feature mining [16, 15, 12]. Time series forecasting requires extrapolation from sequential observations. Language models are designed to discern intricate concepts within temporally correlated sequences, and intuitively appear well-suited for this task. Hence, there exist some preliminary studies that apply LLMs to time series forecasting tasks [10, 17, 19].
However, currently the application of LLMs for time series forecasting is still in its early stage, and the boundaries of this research area are not yet well defined. There are many unexplored problems in this field. For example, existing research lacks exploration into how the performance of LLMs varies when faced with different types of time series inputs. This includes the effectiveness gap for LLMs in predicting data with seasonal and trending patterns versus data without such patterns.
@@
To fill this research gap, in this paper, we focus on LLMs’ preferences for the input time series in time series forecasting under the zero shot prompting setting. Through experiments on both real and synthesized datasets, we find that LLMs perform better in time series with higher trend or seasonal strengths. Our observations also reveal that LLMs perform worse when there are multiple periods within datasets, which may be attributed to the fact that LLMs cannot capture the distinct periods within those datasets. To further discern the LLMs’ preferences for the specific segments of the input data, we design counterfactual experiments involving systematic permutations of input sequences. The findings suggest that LLMs are particularly sensitive to the segment of input sequences closest to the target output.
Based on the above findings, we want to further explore why LLMs forecast well on datasets with higher seasonal strengths. To this end, we require LLMs to tell the period of the datasets through multiple runs. We find that LLMs can mostly recognize the underlying period of a dataset. This can explain the findings of why large language models can forecast time series with high trends or seasonal intensities well, since they can obtain the seasonal pattern inside the datasets.
@@
In light of the above-mentioned findings, we are interested in how to leverage these insights to further improve model performance. To address this, we propose two simple techniques to enhance model performance: incorporating external human knowledge and converting numerical sequences into natural language counterparts. Incorporating supplementary information enables large language models to more effectively grasp the periodic nature of time series data, moving beyond a mere emphasis on the tail of the time series. Transforming numerical data into a natural language format enhances the model’s ability to comprehend and reason, also serving as a beneficial approach. Both approaches improve model performance and contribute to our understanding of LLMs in time series forecasting. The workflow is illustrated in Figure 1.
@@
We investigate the preferences for the input sequences in LLMs in time series forecasting tasks. Our analysis has revealed that LLMs significantly outperform traditional time series forecasting methods without the need for additional fine-tuning. Interestingly, LLMs display superior predictive capabilities when dealing with datasets that have higher trends and seasonal strengths.
We require LLMs to identify the periodicity of datasets across multiple iterations. Our observations indicate that LLMs can effectively recognize the inherent periodic patterns within datasets. This observation answers the question of why LLMs perform well in forecasting time series with higher seasonal strengths, as they can capture the seasonal patterns inherent in the data.
We propose two simple techniques to improve model performance and find that both incorporating external human knowledge into input prompts and paraphrasing input sequences to natural language substantially improve the performance of LLMs in time series forecasting.
@@ 2.1 Large Language Model
We use LLMs as a zero-shot learner for time series forecasting by treating numerical values as text sequences. In this paper, we investigate three close source LLMs, including GPT-3.5-turbo, GPT-4-turbo, and Gemini-1.0-Pro, and one open-source LLMs, i.e., llama-2-13B. The success of LLMs in time series forecasting can significantly depend on correct pre-processing and handling of the data [10]. We followed the pre-processing approach of Gruver [10] and this process involves the following few steps.
Input Pre-processing. In this phase for time series forecasting with LLMs, we perform two pre-processing steps. First, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing. For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to "1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0", introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space. Second, tokenization is equally important, shaping the model’s pattern recognition capabilities. Unlike traditional methods such as byte-pair encoding (BPE) [13], which can disrupt numerical coherence, we use spacing digits which ensures individual tokenization, enhancing pattern discernment. Third, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so that a specific percentile aligns to 1. This facilitates the model’s exposure to varying digit counts and supports the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
@@ 2.2 Time Series Forecasting
Here, represent the -step estimation given the previous -step values . denotes the trained parameters from the model , and denotes the prompt or any other information used for inference. In this paper, we focus predominantly on univariate time series forecasting to investigate the preference and performance of LLMs in univariate time series forecasting under the zero-shot setting.
Motivated by interpretability requirements in real-world scenarios, time series can often be decomposed into the trend component, the seasonal component, and the residual component through the addictive model [5]. The trend component captures the hidden long-term changes in the data, such as the linear or exponential pattern. The seasonal component captures the repeating variation in the data, and the residual component captures the remaining variation in the data after removing the trend and seasonal components. This decomposition offers a method to quantify the properties of time series, which is detailed in subsection 3.2.
Datasets. In this study, we primarily use Darts [11], a benchmark univariate dataset widely recognized in deep learning research, along with many baseline methods. Darts consists of eight real univariate time series datasets, including those with clear patterns, such as the AirPassengerDataset, and irregular datasets, such as the SunspotsDataset. Besides, we employ some other commonly used datasets, such as US Births Dataset[9], TSMC-Stock and Turkeypower datasets [10] and ETT [22] in Sections 5.1 and 5.2 to demonstrate the effectiveness of our proposed methods. A full description of those datasets can be seen in Appendix A.1.
Evaluation Metrics. In this paper, we evaluate model performance with three metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). These metrics are defined as follows:
where denotes the true value, represents the predicted value, and is the sample size.
@@ 3 What are LLMs’ Preferences in Time Series Forecasting?
To explore the preference of LLMs, we first quantify the properties of the input time series to investigate the LLMs’ preferences for time series. Then, to further emphasize our findings, we evaluate the importance of different segments of the input sequence by adding Gaussian noise to the original time series.
@@ 3.1 Analyzing Method
We first compare the performance between LLMs and traditional time series forecasting methods, as shown in Table 9 and Table 10. It is shown that LLMs perform better within most datasets. GPT-4-turbo and Llama-2 perform relatively well on the AirPassengerdataset and the AusBeerdataset with low MAPE. Gemini outperforms GPT-3.5-turbo on time series forecasting and outperforms GPT-4-turbo on some datasets but is on par with GPT-4-turbo overall.
To understand the preferences of the LLMs, we compare our framework using various foundational models, such as GPT-4-turbo and GPT-3.5-turbo, with traditional methods. We also design experiments on synthesized datasets to validate our findings and analyze the impact of the multiple periods. To quantify the LLMs’ preferences towards time series, following [20], we define the strength of the trend and the seasonality as follows:
where , and denote the trend component, the seasonal component and the residual component respectively. The presented indices indicate the trend’s strength and seasonality, providing a measure ranging up to 1. It is easy to find that a higher value indicates a stronger trend or seasonality within the time series. Throughout this paper, we use the word "higher strength" to represent the comparison of the strengths between different datasets. The assessment of strength is not based on a fixed level, as the concepts of "strong" and "weak" vary across different datasets and scenarios.
To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian noise to the original time series to create counterfactual examples. We start by defining a sliding window that constitutes 10% of the total length of the time series, and we set the sliding window to gradually move closer to the output sequence. This method allows us to assess the impact of different segments fairly and thereby infer the interpretability of the time series segments that LLMs predominantly focus on.
3.2 Preferences for Input Sequences
In this subsection, we investigate the input sequence preferences for time series forecasting with LLMs. We conduct experiments on real datasets with GPT-3.5-turbo and GPT-4-turbo, measuring model performance through MAPE. To further validate our findings, we also use GPT-3.5-turbo and Gemini-1.0-Pro to forecast multiple-period time series on synthesized datasets.
| Metrics | GPT4-MAPE | GPT3.5-MAPE | Trend Strength | Seasonal Strength |
| GPT4-MAPE | 1.00000 | 0.987398 | -0.020637 | -0.681440 |
| GPT3.5-MAPE | 0.987398 | 1.00000 | -0.115087 | -0.669983 |
| Trend Strength | -0.020637 | -0.115087 | 1.00000 | 0.508980 |
| Seasonal Strength | -0.681440 | -0.669983 | 0.508980 | 1.00000 |
@@ 3.2.1 Implementation Details
Real Datasets: We conduct experiments on ten real-world datasets, including both those with clear patterns and those with irregular characteristics. The results are shown in Table 6. We apply the Seasonal-Trend decomposition using the LOESS (STL) technique [5] to decompose the original time series into trend, seasonal, and residual components. Subsequently, we compute the strengths of the trend strength and seasonal strength . To further understand the LLMs’ preferences for the specific segments of the input data, we conduct the counterfactual analysis with a systematic permutation to the input time series. We first scale the sequence through max-min normalization. We then define a sliding window that constitutes 10% of the total length of the time series and add Gaussian noise into the data within this window data. Subsequently, the sliding window moves closer to the last known data point.
Refer to subsubsection A.2.1 for detailed information.
Synthesized Datasets: To further validate our findings and investigate the influence of the number of periods on model performance, we generate a dataset using the function . ranges from 0 to 20 and follows the normal distribution . Refer to subsubsection A.2.1 for detailed information.
@@ 3.2.2 Key Findings
After computing the Pearson correlation coefficients (PCC), we observe a nearly strong correlation between the strengths and model performance, showing that LLMs perform better when the input time series has a higher trend and seasonal strength, which is shown in Table 1.In the context of multi-period time series, the model performance worsens as the number of periods increases. It indicates that LLMs may have difficulty recognizing the multiple periods inherent in such datasets. This potentially stems from their inability to adequately capture long-term periods, which can be supported in Figure 8. Besides, for counterfactual analysis, as shown in Figure 2 and Figure 3, there is a noticeable increase in MAPE values when Gaussian noise is added to the latter segments, while the perturbation of the first part of the sequence has little effect on the prediction performance. Our findings reveal that LLMs are more sensitive to the end of input time series when forecasting. We show our full results in Figure 6 and Figure 7. As we move to the right along the x-axis, the closer it gets to the output sequence. We defer the details in Appendix A.
It is also found that the initial part of the sequence has the least impact on the prediction accuracy. For the datasets with high seasonal strengths over 85%, such as WoolyDataset, and MonthlymilkDataset, more than 80% of the length of the time series has almost no effect on the model performance.
4 Why do LLMs Forecast Well on Data with Higher Seasonal Strengths?
Our findings show that LLMs demonstrate enhanced performance in time series forecasting with strong seasonal strengths. This raises the question: Why do LLMs perform well in forecasting datasets with marked seasonal patterns? To explore this phenomenon, we craft prompts that require LLMs to recognize the dataset’s temporal pattern.
This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct seasonal attributes. By explicitly prompting LLMs to predict the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, which sheds light on the mechanisms that underpin their superior performance in such contexts.
@@ 4.1 Implementation Details
To explore the phenomenon that LLMs forecast well on datasets with higher seasonal strengths, we design experiments to verify this phenomenon. We tokenize the input sequence and let the LLMs output the period directly. We use GPT-3.5-turbo, GPT-4-turbo and Gemini-1.0-Pro to predict the periods. We have chosen five datasets with their seasonal strengths exceeding 85%. These datasets are readily available with clear seasonal patterns. In contrast, determining the specific periods of other irregular datasets is challenging, as they have no specific cycles. We record the predicted periods ten times and identify the mode period, which is the most frequently predicted value. We then compare the mode of these ten results with the real period. The mode is selected as the evaluation metric because, when considering the usage characteristics of LLMs, the output of this number best represents the model’s normal performance. We defer our details of the prompt in subsection A.3, and the results are shown in Table 2.
@@ 4.2 Key Findings
According to the results, we find that large language models can mostly determine the periodicity of a dataset. The true periods are determined here by the periodogram, which is commonly used to identify the dominant periods [1]. The multiples of the predicted period also align with the original data cycle. Consequently, we consider the prediction of these multiples to be accurate. We observe that LLMs generally perform well in predicting the period for most datasets with minimal fluctuations. Surprisingly, we discover that in the case of WoolyDataset and AusbeerDataset, which possess relatively short underlying periods, the predicted period is consistently 3 instead of the true period, 4. This discrepancy may be attributed to the LLMs’ tendency to focus on cyclic patterns among individual digits rather than considering the entire sequence as a whole, a phenomenon that could also be interpreted as the model’s identification of the underlying cycle. We leave a comprehensive analysis of this phenomenon in the future.
5 How to Leverage These Insights to Improve the Model’s Performance?
Based on the findings in the previous two sections, our focus is now on how to leverage these findings to further improve model performance. In this paper, we propose two approaches to the user input without additional fine-tuning: for the input prompt, we incorporate additional knowledge of the specific trend and seasonal patterns in the dataset, which gives the model a richer understanding of the underlying patterns. Regarding the input sequence, we transform the time series data into formats resembling natural language sequences rather than relying on the original tokenization. This approach leverages LLMs’ superior capabilities with language sequences. Both methods achieve substantially improved model performance.